OpenStudy (baby456):

4r+5r- 6r= 15+6 Help me i will become fan and give medal to you

OpenStudy (anonymous):

you need to find r?

OpenStudy (baby456):

i don't know i am confused

OpenStudy (anonymous):

4r+5r-6r=15+6 9r-6r=21 3r=21 divide both side by 3. if you divide 3r by 3 ,r will stand alone. then you will divide 21 by 3 r will be = 7

OpenStudy (anonymous):

^^^what she said

OpenStudy (baby456):

thanks you guys

OpenStudy (baby456):

i stick to muy promises

OpenStudy (baby456):

bye